# Efficient Diffusion Models: A Survey

### v2

@@ Abstract
Diffusion models have emerged as powerful generative models capable of producing high-quality contents such as images, videos, audio, 3D, and text demonstrating their potential to revolutionize digital content creation. However, these capabilities come at the cost of their significant computational resources and lengthy generation time, underscoring the critical need to develop efficient techniques for practical deployment. In this survey, we provide a systematic and comprehensive review of research on efficient diffusion models. We organize the literature in a taxonomy consisting of three main categories, covering distinct yet interconnected efficient diffusion model topics from algorithm-level, system-level, and framework perspective, respectively. We have also created a GitHub repository where we organize the papers featured in this survey at https://github.com/AIoT-MLSys-Lab/Efficient-Diffusion-Model-Survey. We hope our survey can serve as a valuable resource to help researchers and practitioners gain a systematic understanding of efficient diffusion model research and inspire them to contribute to this important and exciting field.
@@ 2.1.1 DDPM
$q(\mathbf{x}_{1:T}|\mathbf{x}_{0}):=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),\quad q(\mathbf{x}_{t}|\mathbf{x}_{t-1}):=\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{t-1},\beta_{t}I)$ (1)
$\beta_{t}$ is defined as the variance of the noise added at each timestep. We then convert this to $\alpha_{t}=1-\beta_{t}$. Additionally, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ is defined as the cumulative product of $\alpha_{t}$, following the formulation by Sohl-Dickstein et al. (Sohl-Dickstein et al., 2015). This cumulative product allows for modeling the transition from the original data $\mathbf{x}_{0}$ to $\mathbf{x}_{t}$ as a Gaussian distribution:
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$ (2)
@@
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ (3)
@@
$p_{\theta}(\mathbf{x}_{0:T}):=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}),\quad p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}):=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$ (4)
@@
$L_{\text{simple}}(\theta):=\mathbb{E}_{t,\mathbf{x_{0}},\bm{\epsilon}}\left[\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t)\|^{2}\right]$ (6)
@@ 2.1.2 Score Matching
$L(\bm{\theta})=\frac{1}{2}\mathbb{E}_{p_{d}(\mathbf{x})}\left[\|\bm{s}_{m}(\mathbf{x};\bm{\theta})-\bm{s}_{d}(\mathbf{x})\|^{2}\right]$ (7)
@@
$J(\bm{\theta})=\mathbb{E}_{p_{d}(\mathbf{x})}\left[\text{tr}(\nabla_{\mathbf{x}}\bm{s}_{m}(\mathbf{x};\bm{\theta}))+\frac{1}{2}\|\bm{s}_{m}(\mathbf{x};\bm{\theta})\|^{2}\right]$ (8)
In this expression, $\text{tr}(\cdot)$ denotes the trace of the Hessian matrix of $\bm{s}_{m}(\mathbf{x};\bm{\theta})$. The constant $C$ is independent of $\bm{\theta}$ and can be ignored for optimization purposes. The final form of the unbiased estimator used for training is:
$\hat{J}(\bm{\theta};\mathbf{x}_{1}^{N})=\frac{1}{N}\sum_{i=1}^{N}\left[\text{tr}(\nabla_{\mathbf{x}}\bm{s}_{m}(\mathbf{x}_{i};\bm{\theta}))+\frac{1}{2}\|\bm{s}_{m}(\mathbf{x}_{i};\bm{\theta})\|^{2}\right]$ (9)
@@ 2.1.3 Solvers
$\bar{\mathbf{w}}$ is a standard Wiener process when time flows backward from $T$ to 0, $dt$ is an infinitesimal negative timestep and $\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})$ represent the score function that we mentioned in Eq.(10). In the diffusion process, reverse-time SDE converts noise into data gradually. The complete SDE process is shown in Figure 2.
@@ 2.1.4 Flow Matching
where: $u_{t}(x)$ generates a predefined probability density path $p_{t}(x)$, $p_{0}(x)=\mathcal{N}(x|0,I)$ is a simple prior distribution (e.g., Gaussian noise), and $p_{1}(x)\approx q(x)$ approximates the data distribution.
@@
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,q(x_{1}),p_{t}(x|x_{1})}\|v_{t}(x)-u_{t}(x|x_{1})\|^{2},$ (21)
The second key theorem in the paper proves that the FM and CFM objectives have identical gradients with respect to $\theta$, so optimizing CFM is equivalent to optimizing FM, but computationally simpler. Compared to score matching, flow matching is faster because score matching typically requires computing Hessian matrices in Eq(9), while CFM cleverly avoids such complex calculations by only computing the L2 distance between vector fields. In conclusion, it’s worth noting that Flow Matching can be theoretically connected with other diffusion model formulations. As summarized by Gao et al. (2024), flow matching, score-based models, and traditional diffusion models can be understood within a unified mathematical framework. These approaches essentially represent different parameterizations of the same continuous-time process. While traditional diffusion models like DDPM (Ho et al., 2020) learn to estimate noise or score functions, and score-based models directly optimize gradient fields (Hyvärinen & Dayan, 2005), flow matching learns the vector field itself, directly capturing the transformation map between distributions.
@@ 3.1.2 Loss Formulation
While effective for dimensionality reduction, score estimation still faces challenges in low data density regions where data samples are sparse.
@@
Building upon sliced score matching, to address the issue of inaccurate score estimation in low data density regions, Song & Ermon (2019) introduces a novel generative framework that employs Langevin dynamics to produce samples based on estimated gradients of the data distribution $p_{\text{data}}(\mathbf{x})$. They proposed Noise Conditional Score Networks (NCSN) $s_{\theta}(\mathbf{x},\sigma)$, which jointly estimate scores across multiple noise-perturbed data distributions. By conditioning on a geometric sequence of noise levels $\sigma_{3}>\sigma_{2}>\sigma_{1}$, a single network learns to estimate scores for distributions ranging from highly smoothed $p_{\sigma_{3}}(\mathbf{x})$ that fill low-density regions to concentrated $p_{\sigma_{1}}(\mathbf{x})$ that preserve the structure of the original data manifold. This unified training approach enables robust score estimation across the entire data space. Following a similar derivation, as Song et al. (2021), Dockhorn et al. (2021) introduces Coupled Langevin Dynamics (CLD), redefining the score matching objective within the CLD framework. Unlike traditional score matching methods that inject noise directly into the data space, CLD simplifies the task by only requiring the model to learn the score of the conditional distribution $p_{t}(v_{t}\mid x_{t})$, where noise is injected into an auxiliary variable $v_{t}$ coupled with the data.
Figure 5: Illustration of data-dependent adaptive priors for diffusion processes across different modalities, demonstrating how tailored priors improve generation on quality.
Rectified Flow. As illustrated in Figure 4, Rectified Flow, proposed by (Liu et al., 2022b; Liu, 2022), introduces a method for training ordinary differential equation (ODE) models by learning straight transport paths between two distributions, $\pi_{0}$ and $\pi_{1}$.
The key idea is to minimize the transport cost by ensuring that the learned trajectory between these two distributions follows the most direct route, which can be computationally efficient to simulate. Unlike traditional diffusion models, which may follow roundabout paths, Rectified Flow leverages a simpler optimization problem to create a straight flow, thereby improving both training efficiency and the quality of the generated samples. Flow Matching (Lipman et al., 2022), as discussed in  §2.1.4, introduces the Optimal Transport (OT) approach as a significant advancement in generative modeling. The key innovation of the OT method lies in how it parameterizes conditional probability paths. It represents conditional probability paths in Eq(19) using Gaussian distributions with means $\mu_{t}(x)=tx_{1}$ and standard deviations $\sigma_{t}(x)=1-(1-\sigma_{\min})t$ that change linearly with time. This linear parameterization produces a vector field with constant direction, enabling transitions from noise to data along straight-line trajectories as well.
@@
InstaFlow’s success highlights the potential of Rectified Flow in dramatically reducing the computational cost of diffusion models while maintaining high output quality.
Following InstaFlow, Yan et al. (2024) proposed PeRFlow, further extending the Rectified Flow concept to create a more flexible and universally applicable acceleration method. PeRFlow divides the sampling process into multiple time windows, applying the reflow operation to each interval, creating piecewise linear flows that allow for more nuanced trajectory optimization. Through carefully designed parameterizations, PeRFlow models can inherit knowledge from pretrained diffusion models, achieving fast convergence and superior transfer ability. This approach positions PeRFlow as a universal plug-and-play accelerator compatible with various workflows based on pretrained diffusion models. While Rectified Flow showed great promise, there was still room for improvement, especially in low Number of Function Evaluations (NFE) settings. Addressing this, Lee et al. (2024a) focused on enhancing the training process of Rectified Flows. They discovered that a single iteration of the Reflow algorithm is often sufficient to learn nearly straight trajectories and introduced a U-shaped timestep distribution and LPIPS-Huber premetric to improve one-round training. These improvements led to significant enhancements in FID scores, particularly in low NFE settings, outperforming state-of-the-art distillation methods on various datasets. Most recently, Zhu et al. (2024) proposed SlimFlow, a method designed to address the joint compression of inference steps and model size within the Rectified Flow framework, introducing Annealing Reflow to address initialization mismatches between large teacher models and small student models, and developing Flow-Guided Distillation to improve performance on smaller student models.
@@ 3.1.3 Training Tricks
Lee et al. (2021) introduced PriorGrad, which enhances the training of diffusion models for speech synthesis by using an adaptive prior derived from conditional data statistics. This method significantly improves the efficiency of the denoising process during training, leading to faster convergence while enhancing the perceptual quality and robustness of generated samples, even with smaller network capacities.
Vignac et al. (2022) proposed DiGress, a discrete denoising diffusion model for graph generation. By leveraging data-dependent priors, this model better captures the discrete nature of graph data, improving training efficiency and the quality of generated graphs, particularly for applications like chemical molecular structures and social networks.
@@
Similarly, another method tackles the optimization dilemma in latent diffusion models (LDMs) by introducing extra loss supervision through the VA-VAE framework (Yao et al., 2025). This approach aligns the latent space of a variational autoencoder with a pre-trained vision foundation model, using a loss term to ensure the latent representations capture the data distribution more effectively. This supervision mitigates the trade-off between reconstruction fidelity and generation quality, a common challenge in LDMs where increasing latent dimensionality often hampers generative performance. Combined with an optimized LightningDiT architecture, this method achieves an FID of 1.35 on ImageNet 256x256, with training convergence accelerated by over 21$\times$ compared to baseline DiT models. The extra loss supervision in VA-VAE enhances the model’s ability to balance reconstruction and generation tasks, leading to both efficiency gains and superior sample quality.
@@ 3.2.1 LoRA
The complete process is illustrated in Figure 8. A key advantage of this design lies in its deployment efficiency, where the explicit computation and storage of $W=W_{0}+BA$ enables standard inference procedures without introducing additional latency. Originally proposed for fine-tuning Large Language Models (LLMs), LoRA has demonstrated remarkable parameter efficiency and memory reduction in model adaptation. While predominantly utilized in LLM fine-tuning, recent research has extended its application to diffusion models, indicating its potential as a versatile adaptation technique across different deep learning architectures. LCM-LoRA (Luo et al., 2023b) proposes a universal acceleration approach for diffusion models. As shown in Figure 10, this method achieves fast sampling by adding an Acceleration vector $\tau_{LCM}$ to the Base LDM Rombach et al. (2022). This module adopts LoRA (Hu et al., 2021) to attach low-rank matrices to the original model without architectural modifications. For customized diffusion models that are fine-tuned for specific text-to-image generation tasks, the task-specific LoRA ($\tau^{\prime}$) and acceleration LoRA ($\tau_{LCM}$) can be linearly combined through Eq.(30) to achieve fast inference while maintaining generation quality. More importantly, it provides a plug-and-play solution that reduces sampling steps from dozens to around 4, while maintaining compatibility with any pre-trained text-to-image diffusion model.
$\tau^{\prime}_{LCM}=\lambda_{1}\tau^{\prime}+\lambda_{2}\tau_{LCM}$ (30)
@@
$\epsilon_{\theta^{*}}(x,c_{t},t)\leftarrow\epsilon_{\theta}(x,c_{t},t)+\eta(\epsilon_{\theta}(x,c_{+},t)-\epsilon_{\theta}(x,c_{-},t))$ (31)
Here, $\epsilon_{\theta}$ represents the denoising model’s prediction, and $\eta$ is a guidance scale. With this formulation, the method enables smooth control over concept strength through the guidance scale $\eta$ while maintaining concept independence in the learned directions. By leveraging LoRA’s parameter-efficient nature, it achieves precise attribute manipulation with minimal computational overhead.

### v3

@@ Abstract
Diffusion models have emerged as powerful generative models capable of producing high-quality contents such as images, videos, audio, and text, demonstrating their potential to revolutionize digital content generation. However, these capabilities come at the cost of their significant resource demands and lengthy generation time, underscoring the need to develop efficient techniques for practical deployment. In this survey, we provide a systematic and comprehensive review of research on efficient diffusion models. We organize the literature in a taxonomy consisting of three main categories, covering distinct yet interconnected efficient diffusion model topics from algorithm-level, system-level, and framework perspective, respectively. We have also created a GitHub repository where we organize the papers featured in this survey at https://github.com/AIoT-MLSys-Lab/Efficient-Diffusion-Model-Survey. We hope our survey can serve as a valuable resource to help researchers and practitioners gain a systematic understanding of efficient diffusion model research and inspire them to contribute to this important and exciting field.
@@ 2.1.1 DDPM
$q(\mathbf{x}_{1:T}|\mathbf{x}_{0}):=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),\quad q(\mathbf{x}_{t}|\mathbf{x}_{t-1}):=\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{t-1},\beta_{t}I),$ (1)
where $\beta_{t}$ is defined as the variance of the noise added at each timestep. We then convert this to $\alpha_{t}=1-\beta_{t}$. Additionally, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ is defined as the cumulative product of $\alpha_{t}$, following the formulation by Sohl-Dickstein et al. (Sohl-Dickstein et al., 2015). This cumulative product allows for modeling the transition from the original data $\mathbf{x}_{0}$ to $\mathbf{x}_{t}$ as a Gaussian distribution:
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}).$ (2)
@@
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}.$ (3)
@@
$p_{\theta}(\mathbf{x}_{0:T}):=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}),\quad p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}):=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t)),$ (4)
@@
$L_{\text{simple}}(\theta):=\mathbb{E}_{t,\mathbf{x_{0}},\bm{\epsilon}}\left[\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t)\|^{2}\right].$ (6)
@@ 2.1.2 Score Matching
$L(\bm{\theta})=\frac{1}{2}\mathbb{E}_{p_{d}(\mathbf{x})}\left[\|\bm{s}_{m}(\mathbf{x};\bm{\theta})-\bm{s}_{d}(\mathbf{x})\|^{2}\right].$ (7)
@@
$J(\bm{\theta})=\mathbb{E}_{p_{d}(\mathbf{x})}\left[\text{tr}(\nabla_{\mathbf{x}}\bm{s}_{m}(\mathbf{x};\bm{\theta}))+\frac{1}{2}\|\bm{s}_{m}(\mathbf{x};\bm{\theta})\|^{2}\right],$ (8)
where $\text{tr}(\cdot)$ denotes the trace of the Hessian matrix of $\bm{s}_{m}(\mathbf{x};\bm{\theta})$. The constant $C$ is independent of $\bm{\theta}$ and can be ignored for optimization purposes. The final form of the unbiased estimator used for training is:
$\hat{J}(\bm{\theta};\mathbf{x}_{1}^{N})=\frac{1}{N}\sum_{i=1}^{N}\left[\text{tr}(\nabla_{\mathbf{x}}\bm{s}_{m}(\mathbf{x}_{i};\bm{\theta}))+\frac{1}{2}\|\bm{s}_{m}(\mathbf{x}_{i};\bm{\theta})\|^{2}\right].$ (9)
@@ 2.1.3 Solvers
where $\bar{\mathbf{w}}$ is a standard Wiener process when time flows backward from $T$ to 0, $dt$ is an infinitesimal negative timestep and $\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})$ represent the score function that we mentioned in Eq.(10). In the diffusion process, reverse-time SDE converts noise into data gradually. The complete SDE process is shown in Figure 2.
@@ 2.1.4 Flow Matching
where $u_{t}(x)$ generates a predefined probability density path $p_{t}(x)$, $p_{0}(x)=\mathcal{N}(x|0,I)$ is a simple prior distribution (e.g., Gaussian noise), and $p_{1}(x)\approx q(x)$ approximates the data distribution.
@@
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,q(x_{1}),p_{t}(x|x_{1})}\|v_{t}(x)-u_{t}(x|x_{1})\|^{2}.$ (21)
The second key theorem in the paper proves that the FM and CFM objectives have identical gradients with respect to $\theta$, so optimizing CFM is equivalent to optimizing FM, but computationally simpler. Compared to score matching, flow matching is faster because score matching typically requires computing Hessian matrices in Eq.(9), while CFM cleverly avoids such complex calculations by only computing the L2 distance between vector fields. In conclusion, it is worth noting that Flow Matching can be theoretically connected with other diffusion model formulations. As summarized by Gao et al. (2024), flow matching, score-based models, and traditional diffusion models can be understood within a unified mathematical framework. These approaches essentially represent different parameterizations of the same continuous-time process. While traditional diffusion models such as DDPM (Ho et al., 2020) learn to estimate noise or score functions, and score-based models directly optimize gradient fields (Hyvärinen & Dayan, 2005), flow matching learns the vector field itself, directly capturing the transformation map between distributions.
@@ 3.1.2 Loss Formulation
While effective for dimensionality reduction, score estimation still faces challenges in low data density regions where data samples are sparse. Building upon sliced score matching, to address the issue of inaccurate score estimation in low data density regions, Song & Ermon (2019) introduces a novel generative framework that employs Langevin dynamics to produce samples based on estimated gradients of the data distribution $p_{\text{data}}(\mathbf{x})$. They proposed Noise Conditional Score Networks (NCSN) $s_{\theta}(\mathbf{x},\sigma)$, which jointly estimate scores across multiple noise-perturbed data distributions. By conditioning on a geometric sequence of noise levels $\sigma_{3}>\sigma_{2}>\sigma_{1}$, a single network learns to estimate scores for distributions ranging from highly smoothed $p_{\sigma_{3}}(\mathbf{x})$ that fill low-density regions to concentrated $p_{\sigma_{1}}(\mathbf{x})$ that preserve the structure of the original data manifold. This unified training approach enables robust score estimation across the entire data space. Following a similar derivation, as Song et al. (2021), Dockhorn et al. (2021) introduces Coupled Langevin Dynamics (CLD), redefining the score matching objective within the CLD framework. Unlike traditional score matching methods that inject noise directly into the data space, CLD simplifies the task by only requiring the model to learn the score of the conditional distribution $p_{t}(v_{t}\mid x_{t})$, where noise is injected into an auxiliary variable $v_{t}$ coupled with the data.
@@
Rectified Flow. As illustrated in Figure 4, Rectified Flow, proposed by (Liu et al., 2022b; Liu, 2022), introduces a method for training ordinary differential equation (ODE) models by learning straight transport paths between two distributions, $\pi_{0}$ and $\pi_{1}$. The key idea is to minimize the transport cost by ensuring that the learned trajectory between these two distributions follows the most direct route, which can be computationally efficient to simulate. Unlike traditional diffusion models, which may follow roundabout paths, Rectified Flow leverages a simpler optimization problem to create a straight flow, thereby improving both training efficiency and the quality of the generated samples. Flow Matching (Lipman et al., 2022), as discussed in §2.1.4, introduces the Optimal Transport (OT) approach as a significant advancement in generative modeling. The key innovation of the OT method lies in how it parameterizes conditional probability paths. It represents conditional probability paths in Eq.(19) using Gaussian distributions with means $\mu_{t}(x)=tx_{1}$ and standard deviations $\sigma_{t}(x)=1-(1-\sigma_{\min})t$ that change linearly with time. This linear parameterization produces a vector field with constant direction, enabling transitions from noise to data along straight-line trajectories as well.
@@
InstaFlow’s success highlights the potential of Rectified Flow in dramatically reducing the computational cost of diffusion models while maintaining high output quality. Following InstaFlow, Yan et al. (2024) proposed PeRFlow, further extending the Rectified Flow concept to create a more flexible and universally applicable acceleration method. PeRFlow divides the sampling process into multiple time windows, applying the reflow operation to each interval, creating piecewise linear flows that allow for more nuanced trajectory optimization. Through carefully designed parameterizations, PeRFlow models can inherit knowledge from pretrained diffusion models, achieving fast convergence and superior transfer ability. This approach positions PeRFlow as a universal plug-and-play accelerator compatible with various workflows based on pretrained diffusion models. While Rectified Flow showed great promise, there was still room for improvement, especially in low Number of Function Evaluations (NFE) settings. Addressing this, Lee et al. (2024a) focused on enhancing the training process of Rectified Flows. They discovered that a single iteration of the Reflow algorithm is often sufficient to learn nearly straight trajectories and introduced a U-shaped timestep distribution and LPIPS-Huber premetric to improve one-round training. These improvements led to significant enhancements in FID scores, particularly in low NFE settings, outperforming state-of-the-art distillation methods on various datasets. Most recently, Zhu et al. (2024) proposed SlimFlow, a method designed to address the joint compression of inference steps and model size within the Rectified Flow framework, introducing Annealing Reflow to address initialization mismatches between large teacher models and small student models, and developing Flow-Guided Distillation to improve performance on smaller student models.
@@ 3.1.3 Training Tricks
Figure 5: Illustration of data-dependent adaptive priors for diffusion processes across different modalities.
@@
Lee et al. (2021) introduced PriorGrad, which enhances the training of diffusion models for speech synthesis by using an adaptive prior derived from conditional data statistics. This method significantly improves the efficiency of the denoising process during training, leading to faster convergence while enhancing the perceptual quality and robustness of generated samples, even with smaller network capacities. Vignac et al. (2022) proposed DiGress, a discrete denoising diffusion model for graph generation. By leveraging data-dependent priors, this model better captures the discrete nature of graph data, improving training efficiency and the quality of generated graphs, particularly for applications like chemical molecular structures and social networks.
@@
Similarly, another method tackles the optimization dilemma in latent diffusion models (LDMs) by introducing extra loss supervision through VA-VAE (Yao et al., 2025). This approach aligns the latent space of a variational autoencoder with a pre-trained vision foundation model, using a loss term to ensure the latent representations capture the data distribution more effectively. This supervision mitigates the trade-off between reconstruction fidelity and generation quality, a common challenge in LDMs where increasing latent dimensionality often hampers generative performance. Combined with an optimized LightningDiT architecture, it achieves an FID of 1.35 on ImageNet 256x256, with training convergence accelerated by over 21$\times$ compared to baseline DiT models. The extra loss supervision in VA-VAE enhances the model’s ability to balance reconstruction and generation tasks, leading to both efficiency gains and superior sample quality.
@@ 3.2.1 LoRA
The complete process is illustrated in Figure 9. A key advantage of this design lies in its deployment efficiency, where the explicit computation and storage of $W=W_{0}+BA$ enables standard inference procedures without introducing additional latency. Originally proposed for fine-tuning Large Language Models (LLMs), LoRA has demonstrated remarkable parameter efficiency and memory reduction in model adaptation. While predominantly utilized in LLM fine-tuning, recent research has extended its application to diffusion models, indicating its potential as a versatile adaptation technique across different deep learning architectures.
LCM-LoRA (Luo et al., 2023b) proposes a universal acceleration approach for diffusion models. As shown in Figure 10, this method achieves fast sampling by adding an Acceleration vector $\tau_{LCM}$ to the Base LDM Rombach et al. (2022). This module adopts LoRA (Hu et al., 2021) to attach low-rank matrices to the original model without architectural modifications. For customized diffusion models that are fine-tuned for specific text-to-image generation tasks, the task-specific LoRA ($\tau^{\prime}$) and acceleration LoRA ($\tau_{LCM}$) can be linearly combined through Eq.(30) to achieve fast inference while maintaining generation quality. More importantly, it provides a plug-and-play solution that reduces sampling steps from dozens to around 4, while maintaining compatibility with any pre-trained text-to-image diffusion model.
$\tau^{\prime}_{LCM}=\lambda_{1}\tau^{\prime}+\lambda_{2}\tau_{LCM}.$ (30)
@@
$\epsilon_{\theta^{*}}(x,c_{t},t)\leftarrow\epsilon_{\theta}(x,c_{t},t)+\eta(\epsilon_{\theta}(x,c_{+},t)-\epsilon_{\theta}(x,c_{-},t)),$ (31)
where $\epsilon_{\theta}$ represents the denoising model’s prediction, and $\eta$ is a guidance scale. With this formulation, the method enables smooth control over concept strength through the guidance scale $\eta$ while maintaining concept independence in the learned directions. By leveraging LoRA’s parameter-efficient nature, it achieves precise attribute manipulation with minimal computational overhead.